# Geodesic Trees and Exceptional Directions in FPP on Hyperbolic Groups

### v1

@@ 5.3. Relationships between forward and backward trees
Let $o$ be a base-point in ${\mathbf{H}}^{2}$ as above, $z\in\partial{\mathbf{H}}^{2}$, and $[o,z)$ be a geodesic ray. Let $z_{n}\in[o,z)$ be such that $d(z_{n},o)=n$. Then the balls $B_{n}(z_{n})$ of radius $n$ about $z_{n}$ converge to the unique horodisk ${\mathbb{H}}(z)$ based at $z$ and containing $o$ as a point on its boundary horocycle $\partial{\mathbb{H}}(z)$. This geometric construction generalizes to hyperbolic groups in a straightforward manner by replacing ${\mathbf{H}}^{2}$ with a Cayley graph $\Gamma$, the base-point $o$ by $1$ and $z\in\partial{\mathbf{H}}^{2}$ by $z\in{\partial G}$. More generally, this construction shows that horocycles (i.e. orbits of $N$ in ${\mathbf{H}}^{2}$) are limits of boundaries $\partial D(n)$ of a suitably chosen sequence of $n-$disks where $\partial D(n)$ passes through a fixed point (the origin, say) in ${\mathbf{H}}^{2}$. It is this fact that allows us to move from the first to the second descriptions of ${\mathbf{H}}^{2}$ to deduce ergodic theoretic results (see [EM93] and subsequent works for instance). In our present random discrete setup, it shows up in similar asymptotic behavior of forward and backward trees. There are, however, subtle differences that this analogy does not capture. Question 5.6 below captures one such feature.
@@
In planar (Euclidean) exactly solvable LPP, it was shown in [JRAS23, Theorem 3.9] that exceptional directions are independent of the base-point $g$. In conjunction with the resolution of the N3G problem [Cou11] for planar (Euclidean) exactly solvable LPP, the second part of Question 5.6 above automatically has a positive answer in this case. In the deterministic case, where $\rho$ has an atom at 1, Question 5.6 specializes to the full geodesic language (without an ordering on the generating set) generated by the Cannon automaton.
An approach to resolving Question 5.6 is to use the fact, already stated, that for the backward tree $T(\xi,\omega)$, exceptional directions and their multiplicities are tautologically independent of the base-point $g$. Recall the well-known geometric limiting construction of disks to horodisks describes earlier in this subsection. We then ask the following.
Question 5.7.
@@
The strongest sense in which such a convergence could occur is Gromov-Hausdorff. We do not expect a positive answer in such a strong sense. However, one might expect a positive answer if one wants only a convergence in measure. This is due to the following fact: if $z_{n}$ is chosen uniformly randomly from the ball $B_{n}(1)$ of radius $n$, then the uniform probability on $B_{n}(1)$ converges to the Patterson-Sullivan measure on ${\partial G}$. Hence, an appropriate modification of Benjamini-Schramm convergence might yield a positive answer to Question 5.7.
@@ 5.4. Generalizations
3c) Theorem 3.21 uses only basic properties of topological dimension. Hence, it goes through for $X$ satisfying Assumption 5.8, and 5.9.
3d) For the same reasons as (3c) above, Theorem 3.27 goes through for $X$ satisfying Assumption 5.8, and 5.9.
@@
Thus Theorem 4.1 goes through for $X$ satisfying Assumptions 5.8 and 5.10.

### v2

@@ 5.4. Relationships between forward and backward trees
Let $o$ be a base-point in ${\mathbf{H}}^{2}$ as above, $z\in\partial{\mathbf{H}}^{2}$, and $[o,z)$ be a geodesic ray. Let $z_{n}\in[o,z)$ be such that $d(z_{n},o)=n$. Then the balls $B_{n}(z_{n})$ of radius $n$ about $z_{n}$ converge to the unique horodisk ${\mathbb{H}}(z)$ based at $z$ and containing $o$ as a point on its boundary horocycle $\partial{\mathbb{H}}(z)$. This geometric construction generalizes to hyperbolic groups in a straightforward manner by replacing ${\mathbf{H}}^{2}$ with a Cayley graph $\Gamma$, the base-point $o$ by $1$ and $z\in\partial{\mathbf{H}}^{2}$ by $z\in{\partial G}$. More generally, this construction shows that horocycles (i.e. orbits of $N$ in ${\mathbf{H}}^{2}$) are limits of boundaries $\partial D(n)$ of a suitably chosen sequence of $n-$disks where $\partial D(n)$ passes through a fixed point (the origin, say) in ${\mathbf{H}}^{2}$. It is this fact that allows us to move from the first to the second descriptions of ${\mathbf{H}}^{2}$ to deduce ergodic theoretic results (see [EM93] and subsequent works for instance). In our present random discrete setup, it shows up in similar asymptotic behavior of forward and backward trees. There are, however, subtle differences that this analogy does not capture. Question 5.11 below captures one such feature.
@@
In planar (Euclidean) exactly solvable LPP, it was shown in [JRAS23, Theorem 3.9] that exceptional directions are independent of the base-point $g$. In conjunction with the resolution of the N3G problem [Cou11] for planar (Euclidean) exactly solvable LPP, the second part of Question 5.11 above automatically has a positive answer in this case. In the deterministic case, where $\rho$ has an atom at 1, Question 5.11 specializes to the full geodesic language (without an ordering on the generating set) generated by the Cannon automaton.
An approach to resolving Question 5.11 is to use the fact, already stated, that for the backward tree $T(\xi,\omega)$, exceptional directions and their multiplicities are tautologically independent of the base-point $g$. Recall the well-known geometric limiting construction of disks to horodisks describes earlier in this subsection. We then ask the following.
Question 5.12.
@@
The strongest sense in which such a convergence could occur is Gromov-Hausdorff. We do not expect a positive answer in such a strong sense. However, one might expect a positive answer if one wants only a convergence in measure. This is due to the following fact: if $z_{n}$ is chosen uniformly randomly from the ball $B_{n}(1)$ of radius $n$, then the uniform probability on $B_{n}(1)$ converges to the Patterson-Sullivan measure on ${\partial G}$. Hence, an appropriate modification of Benjamini-Schramm convergence might yield a positive answer to Question 5.12.
@@ 5.5. Generalizations
3c) Theorem 3.21 uses only basic properties of topological dimension. Hence, it goes through for $X$ satisfying Assumption 5.13, and 5.14.
3d) For the same reasons as (3c) above, Theorem 3.27 goes through for $X$ satisfying Assumption 5.13, and 5.14.
@@
Thus Theorem 4.1 goes through for $X$ satisfying Assumptions 5.13 and 5.15.